OpenStudy (anonymous):

The numerator and denominator of a fraction are in the ratio of 3 to 5. If the numerator and denominator are both decreased by 2, the fraction is now equal to . If n = the numerator and d = the denominator, which of the following systems of equations could be used to solve the problem? 5n = 3d and n - 2 = 2d - 4 5n = 3d and 2n - 4 = d - 2 3n = 5d and 2n - 4 = d - 2

OpenStudy (anonymous):

we will never solve this problem since u miss a part

OpenStudy (anonymous):

the fraction is now equal to .: ???????????????????????????????

OpenStudy (anonymous):

my bad its \[\frac{ 1 }{ 2}\]

OpenStudy (anonymous):

isnt it the second one?

OpenStudy (anonymous):

for the first equation we have (n/d)=(3/5) the second equation is ( (n-2)/(d-2) ) = 1/2